OpenStudy (anonymous):

Pre-Cal show that the given function is one-to-one and nd its inverse. Check your answers algebraically. Verify that the range of f is the domain of f^-1 and vice-versa. f(x) =(x-2/3)+4

OpenStudy (anonymous):

f[f^-1(x)]=((x-2/3) - 2/3)+4) this is where i get stuck. help. how do i distribute that x-2/3

OpenStudy (cwrw238):

i'm a little confused with this - in the formula for f(x) is this x - two thirds or (x-2) / 3 ?

OpenStudy (anonymous):

f(x) =(x-2/3)+4 but i have to figure it out for f^-1(x) and f(x)=f^-1(x)

OpenStudy (anonymous):

@cwrw238

OpenStudy (cwrw238):

well first simplify f(x): f(x) = x - 2/3 + 12/3 = x + 10/3 x = f(x) - 10/3 giving the inverse f-^1(x) = x - 10/3

OpenStudy (cwrw238):

to check that this is correct inverse f(f-1^x)) should equal x

OpenStudy (anonymous):